# CSEval: Towards Automated, Multi-Dimensional, and Reference-Free Counterspeech Evaluation using Auto-Calibrated LLMs

### v1

@@ 6 Results
We adopt the same approach suggested by Zhong et al. (2022) and Liu et al. (2023a) to evaluate different CS generation metrics using CS-level Spearman ($\rho$) and Kendall-Tau ($\tau$) correlation. We report the correlation scores between automated metrics and human judgments in Table 1. The first part of Table 1 shows metrics that compare lexical or semantic similarity between the model output and reference CS. We find that overlap-based metrics perform poorly on almost all dimensions, with metrics such as BLEU, ROUGE, and METEOR even displaying negative correlations in some cases. This observation is consistent with similar studies in summarization and dialogue generation, where overlap-based metrics are shown to be poor multidimensional evaluators Fabbri et al. (2021); Zhong et al. (2022); Liu et al. (2023a). The second part shows results of metrics that use neural networks to learn from human ratings of text quality. BARTScore particularly achieves better correlations than the similarity-based metrics, suggesting that it is more reliable for CS evaluation.
@@
Lastly, we observe that LLM-based evaluators achieve the highest correlations with human judgements across all dimensions, underlying their reliability as reference-free counterspeech evaluators. Our proposed ACE method with GPT-4 as the base model consistently outperforms all the other LLM baselines, especially for relevance, coherence, and suitableness. Furthermore, we observe that for each of the three LLMs, prompting methods with CoT (G-EVAL and ACE) perform much better than their respective zero-shot counterparts. This validates our hypothesis that CoT in the form of evaluation steps helps LLMs achieve improved alignment, bringing them closer towards human reasoning for counterspeech evaluation.
@@ 7 Analysis
In Table 1, we compare the performance of ACE with and without auto-calibrated CoT on the CSEval benchmark. shows that ACE (GPT-4) outperforms G-EVAL (GPT-4) on both Spearman and Kendall-Tau correlations. We observe a similar trend for Mistral and Llama models. This suggests that auto-calibration helps align the CoT towards human judgement.
@@
Our proposed framework is designed to independently evaluate a CS across different quality dimensions. This is in line with the broader counterspeech literature, which supports the notion that there is no single, absolute metric or aspect that completely represents a counterspeech’s quality or effectiveness Benesch (2014); Chung et al. (2024). However, research following CS generation often requires to relatively compare different NLG methods. To aid this, we experiment with the efficacy of having a unified score for CS evaluation. Specifically, for a given model output, we compute the unified score as a mean of individual scores – relevance, coherence, aggressiveness, and suitableness. For this experiment, we construct a meta-evaluation by randomly sampling model outputs from CSEval. For a given HS, we ask human evaluators to rank model outputs from best to worst according to their preference444 We conduct the meta-evaluation on 372 randomly selected data points from the CSEval test set, representing approximately 15% of the total set. Each data point was independently ranked by three human annotators, and the final rankings were determined by averaging the individual rankings from the three annotators.. Table 2 shows the system-level ranking performance of Llama, Mistral, and GPT-4, computed using the Normalized Discounted Cumulative Gain (NDCG) score. We observe that all LLM-based methods show excellent ranking performance, displaying alignment with human preferences. These results support the use of a unified score, indicating that it can be effective in comparing the relative quality of CS generation models.
@@
LLama-3 (ACE) $0.890$
Mistral (ACE) $0.891$
GPT-4 (ACE) $0.898$
@@ 8 Conclusion
This paper presents CSEval, a novel multidimensional framework for evaluating the quality of automated counterspeech against online hate speech. Our work addresses a critical gap in the current research landscape of automated counterspeech generation by providing a comprehensive and automated approach for assessing counterspeech quality along four key dimensions: context relevance, aggressiveness, argument coherence, and suitableness. We observe that traditional similarity-based evaluation metrics, while prevalent, often fail to capture the nuanced complexity of effective counterspeech. We further introduce ACE, a prompt-based evaluation method with an auto-calibrated chain-of-thoughts mechanism, which leverages LLMs to offer a more refined and human-aligned evaluation. Our experiments with multiple automated metrics on the CSEval dataset illustrate that ACE displays a significant improvement in correlation with human judgment, particularly when juxtaposed against existing evaluation methods. In conclusion, while our framework marks a significant advance in the automated evaluation of counterspeech, it also underscores the complexity and multi-faceted nature of this domain.
@@ 9 Limitations
Our evaluation metrics measure the quality and effectiveness of counterspeeches in a controlled setting, using expert annotations. However, the real-world impact of counterspeeches on social media platforms depends on various factors, such as the reach of content, the audience reactions, and the interactions between hateful and counter speech. These aspects are not captured by our metrics, and are beyond the scope of this paper. Future work could investigate the effectiveness of counterspeeches in the wild, using online experiments or longitudinal studies. Furthermore, our study focuses on four dimensions of text quality that are specific to counterspeech generation: relevance, coherence, aggressiveness, and appropriateness. However, there are other general dimensions of text quality, such as fluency, naturalness, and engagingness, that could also be relevant for evaluating counterspeeches. We do not include these dimensions in our study, as resources to evaluate such quality aspects of generated text have been extensively studied and made available through other open-source datasets like summarisation Fabbri et al. (2021) and dialog-generation Gopalakrishnan et al. (2023). Future work could explore the trade-offs between different text quality dimensions, and how they affect the perception and reception of counterspeeches.
@@ 10 Ethics Statement
We acknowledge the sensitivity involved in dealing with online hate speech and understand that conducting research in this area may need adherence to strict ethical considerations. First, our dataset is based on a publicly available, open-source dataset in the counterspeech domain. As we were largely dealing with online hate speech in the form of social media posts, we ensure that they were fully anonymised and untraceable to the source user. During the data annotation process, we ensured that each of the annotator was fully aware, and had context about the nature and degree of offensive statements that they were responsible to annotate.

### v2

@@ 6 Results
We adopt the same approach suggested by Zhong et al. (2022) and Liu et al. (2023a) to evaluate different CS generation metrics using CS-level Spearman ($\rho$) and Kendall-Tau ($\tau$) correlation. We report the correlation scores between automated metrics and human judgments in Table 1. The first part of Table 1 shows metrics that compare lexical or semantic similarity between the model output and reference CS. We find that overlap-based metrics perform poorly on almost all dimensions, with metrics such as BLEU, ROUGE, and METEOR even displaying negative correlations in some cases. This observation is consistent with similar studies in summarisation and dialogue generation, where overlap-based metrics are shown to be poor multidimensional evaluators Fabbri et al. (2021); Zhong et al. (2022); Liu et al. (2023a). The second part shows results of metrics that use neural networks to learn from human ratings of text quality. In particular, BERTScore and BARTScore show a higher correlation with human judgement than all other similarity-based metrics. This shows that they are more reliable for CS evaluation.
@@
Lastly, we observe that LLM-based evaluators achieve the highest correlations with human judgements across all dimensions, underlying their reliability as reference-free counterspeech evaluators. Our proposed Auto-CSEval method with GPT-4 as the base model consistently outperforms all the other LLM baselines, especially for relevance, coherence, and suitableness. Furthermore, we observe that for each of the three LLMs, prompting methods with CoT (G-EVAL and Auto-CSEval) perform much better than their respective zero-shot counterparts. This validates our hypothesis that CoT in the form of evaluation steps helps LLMs achieve improved alignment, bringing them closer towards human reasoning for counterspeech evaluation.
@@ 7 Analysis
In Table 1, we compare the performance of Auto-CSEval with and without auto-calibrated CoT on the CSEval benchmark. shows that Auto-CSEval (GPT-4) outperforms G-EVAL (GPT-4) on both Spearman and Kendall-Tau correlations. We observe a similar trend for Mistral and Llama models. This suggests that auto-calibration helps align the CoT towards human judgement.
@@
Our proposed framework is designed to independently evaluate a CS across different quality dimensions. This is in line with the broader counterspeech literature, which supports the notion that there is no single, absolute metric or aspect that completely represents a counterspeech’s quality or effectiveness Benesch (2014); Chung et al. (2024). However, research following CS generation often requires to relatively compare different NLG methods. To aid this, we experiment with the efficacy of having a unified score for CS evaluation. Specifically, for a given model output, we compute the unified score as a mean of individual scores – relevance, coherence, aggressiveness, and suitableness.
Mean $=\frac{\text{Relevance}+(6-\text{Aggressiveness})+\text{Coherence}}{4}$
$\quad+\frac{\left(\frac{\text{Suitableness}-1}{2}\times 4+1\right)}{4}$ (2)
As shown in the equation above, Aggressiveness score is normalised as $6-\text{Aggressiveness}$ since it is a "lower the better" score. Similarly, Suitableness is transformed from the original scale of 1-3 to 1-5. This is the same logic used to compute scores in Figure 1.
For this experiment, we construct a meta-evaluation by randomly sampling model outputs from CSEval. For a given HS, we ask human evaluators to rank model outputs from best to worst according to their preference444 We conduct the meta-evaluation on 372 randomly selected data points from the CSEval test set, representing approximately 15% of the total set. Each data point was independently ranked by three human annotators, and the final rankings were determined by averaging the individual rankings from the three annotators.. Table 2 shows the system-level ranking performance of Llama, Mistral, and GPT-4, computed using the Normalized Discounted Cumulative Gain (NDCG) score. We observe that all LLM-based methods show excellent ranking performance, displaying alignment with human preferences. These results support the use of a unified score, indicating that it can be effective in comparing the relative quality of CS generation models.
@@
LLama-3 (Auto-CSEval) $0.890$
Mistral (Auto-CSEval) $0.891$
GPT-4 (Auto-CSEval) $0.898$
@@ 8 Conclusion
This paper presents CSEval, a novel multidimensional framework for evaluating the quality of automated counterspeech against online hate speech. Our work addresses a critical gap in the current research landscape of automated counterspeech generation by providing a comprehensive and automated approach for assessing counterspeech quality along four key dimensions: context relevance, aggressiveness, argument coherence, and suitableness. We observe that traditional similarity-based evaluation metrics, while prevalent, often fail to capture the nuanced complexity of effective counterspeech. We further introduce Auto-CSEval, a prompt-based evaluation method with an auto-calibrated chain-of-thoughts mechanism, which leverages LLMs to offer a more refined and human-aligned evaluation. Our experiments with multiple automated metrics on the CSEval dataset illustrate that Auto-CSEval displays a significant improvement in correlation with human judgment, particularly when juxtaposed against existing evaluation methods. In conclusion, while our framework marks a significant advance in the automated evaluation of counterspeech, it also underscores the complexity and multi-faceted nature of this domain.
@@ 9 Limitations
Our study focuses on four dimensions of text quality that are specific to counterspeech generation: relevance, coherence, aggressiveness, and appropriateness. However, there are other quality aspects, such as fluency, naturalness, opposition (stance), and specificity Chung et al. (2024); Jones et al. (2024). In future work, we plan to extend CSEval to incorporate these quality dimensions, since it can add value to the notion of "unified score" as discussed in section 7. Recent studies have shown that using LLMs as reference-free evaluators (LLMs-as-a-judge) may lead to unseen problems like bias, prompt sensitivity, leniency, and questionable reasoning Thakur et al. (2025); Li et al. (2025). Furthermore, LLM evaluations may also be affected by decoding strategies (temperature or top-k sampling). Our study does not include any analysis on such weakness of LLMs-as-a-judge. Lastly, we would also like to explore the trade-offs between different text quality dimensions and how they affect the perception and reception of counterspeeches. We leave this to future work.
@@ 10 Ethics Statement
We acknowledge that we are dealing with a sensitive topic of research as it deals with online hate speech. We understand that we must adhere to strict ethical considerations while dealing with hatespeech-related data. First, our dataset is based on a publicly available, open-source dataset in the counterspeech domain. As we were largely dealing with online hate speech in the form of social media posts, we ensure that they were fully anonymised and untraceable to the source user. During the data annotation process, we ensured that each of the annotators was fully aware and had context about the nature and degree of offensive statements that they were responsible to annotate.
@@ C.1 Select of Evaluation Dimensions
Appendix D Statistical Tests
Model Relevance Score Aggressiveness Score Coherence Score Suitableness Score
Mean Std Mean Std Mean Std Mean Std
BLEU_1 -0.007 0.022 -0.147 0.043 0.115 0.040 0.005 0.025
ROUGE_l 0.213 0.037 -0.167 0.020 0.354 0.027 0.200 0.027
METEOR_score 0.178 0.032 -0.207 0.020 0.355 0.025 0.169 0.023
BERT_score 0.227 0.027 -0.077 0.027 0.342 0.020 0.249 0.014
Toxicity -0.019 0.030 0.272 0.018 -0.184 0.037 -0.107 0.028
BART_score 0.272 0.024 -0.198 0.023 0.394 0.024 0.249 0.021
GPT-4 (zeroshot) 0.529 0.031 0.336 0.061 0.444 0.025 0.422 0.039
GPT-4 (GEval) 0.629 0.012 0.392 0.030 0.578 0.031 0.522 0.014
GPT-4 (Auto-CSEval) 0.687 0.010 0.461 0.035 0.597 0.032 0.566 0.014
Table 5: Spearman ($\rho$) correlation results for selected methods based on statistical testing.
Model Relevance Score Aggressiveness Score Coherence Score Suitableness Score
Mean Mean Mean Mean
Auto-CSEval - Best Traditional Metric 0.415 0.189 0.202 0.317
Auto-CSEval - Best Method 0.058 0.069 0.019 0.044
Table 6: Delta values comparing Auto-CSEval against the best traditional metric and best method across evaluation aspects.
To verify the reliability of the correlation scores reported in Table 1, we conduct statistical tests for Spearman ($\rho$) correlation values by averaging results across multiple trials. The details of our experimental setup are as follows:
  • Table 1 reports sample-level Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations of different metrics on the CSEval test set. These results are averaged over the entire dataset.
  • We define a “trial” using two variables: sample size ($n$) and random state ($s$). In each trial, we randomly sample $n$ data points from the CSEval test set using the specified random state $s$.
  • We vary the sample sizes linearly, starting from 100 samples and increasing by 500 samples at each step, i.e., 100, 600, 1100, and so on, up to the size of the test set (13 values in total).
  • Following this, we compute correlation scores (same as Table 1) on the randomly sampled subset ($n$ data points) of CSEval.
  • We repeat this procedure for three different seed values: [1, 2, 3].
Thus, we conduct a total of 39 unique trials (13 sample sizes, $\times$ 3 random states). We then average the scores across all 39 trials to finalise our results. The spearman ($\rho$) correlation results for selected methods are presented in Table 5.
In Table 6, we calculated the delta values between (i) Auto-CSEval versus the best traditional metric and (ii) Auto-CSEval versus the best method. These delta values provide a statistically more reliable comparison than those reported in Table 1 of the original manuscript, as the results are averaged across 13 sample sizes and three seed values. We find that our proposed method, Auto-CSEval, shows an average improvement of 0.190 points over the best traditional metric. This improvement is consistent across four evaluation aspects: relevance, coherence, aggressiveness, and suitableness.